# Taxonomy of differentiationP. Komarov, F. van Breugel, J. N. Kutz

### v1

@@ Abstract
Differentiation is a cornerstone of computing and data analysis in every discipline of science and engineering. Indeed, most fundamental physics laws are expressed as relationships between derivatives in space and time. However, derivatives are rarely directly measurable and must instead be computed, often from noisy, potentially corrupt data streams. There is a rich and broad literature of computational differentiation algorithms, but many impose extra constraints to work correctly, e.g. periodic boundary conditions, or are compromised in the presence of noise and corruption. It can therefore be challenging to select the method best-suited to any particular problem. Here, we review a broad range of numerical methods for calculating derivatives, present important contextual considerations and choice points, compare relative advantages, and provide basic theory for each algorithm in order to assist users with the mathematical underpinnings. This serves as a practical guide to help scientists and engineers match methods to application domains. We also provide an open-source Python package, PyNumDiff, which contains a broad suite of methods for differentiating noisy data.
@@ 1 Introduction
Numerical algorithms including Finite Differences, Spectral Methods, and Finite Elements or Volumes are now mature technologies, with the latter powering commercialized (e.g. COMSOL [comsol], Ansys and Abaqus) and open-source (e.g. Fenics, Firedrake, openFOAM) PDE solvers. Likewise, Automatic Differentiation (AutoDiff) dominates in Machine Learning [jax2018github]. But as practitioners in the rapidly-growing field of data-driven science, we also take particular interest in derivative estimation for noisy data streams. Motivated by the definition of the derivative from calculus, beginners in numerical computation often use finite differences by default, but this turns out to be one of the worst choices for real data. We attempt to give comprehensive coverage to all these schemes, from the narrowly focused to the most general-purpose. The overall trend is that more constrictive situations allow for better performance, so we suggest the use of specialized tools if possible. Or, if faced with a dearth of simplifying assumptions, we have made it easy for the reader to skip ahead to the pertinent category.
@@ 2 Choosing a Method
Figure 2: Suggested procedure for choosing a differentiation method. Methods are boxed with blue, violet, purple, or red to give an overall impression of how computationally efficient, accurate, robust, and user-friendly they are, with blue meaning a method excels in all respects for its specialization and red meaning a solution is hopeless. Groupings in the lowermost node represent many methods, expounded in section 7. Methods and a couple other labels are hyperlinked for quick navigation to relevant sections.
@@
A digest of these with some preferred algorithms is given in Figure 1. Figure 2 shows a more complete and detailed overview of key considerations and methods, organized as a flowchart. The decisions therein are important because different techniques make different assumptions. For example, smoothing methods assume the signal is continuously differentiable up to some order (possibly $C^{\infty}$), and the Kalman filter requires a system model. Under changing assumptions, the most suitable approach changes, so it is important to understand whether data is periodic, noisy, etc. Notably missing among these considerations is dimensionality: Because differentiation is a linear operator, all methods extend equally well to the multidimensional case, either by multivariate chain rule, by using higher-dimensional meshes and basis functions, or by sequential application along data dimensions (in any order). We therefore primarily treat the univariate case in this review, without loss of generality.
@@ 2.1 Recommendations
Down the “noisy signal, smooth derivative” path, the Fourier spectral method (i.e. transform, zero out higher modes, and inverse transform) is ideal for isolating high-frequency noise (section 5), but it should only be used when signals are periodic (can be joined smoothly end-to-end), lest Gibbs phenomenon (Figure 7) corrupt the approximation. Otherwise, one should exploit a model of system dynamics and noise when possible (section 6), although system models can be challenging to derive beyond simple systems and can be of limited utility unless fairly accurate. In the naive case (section 7), sophisticated methods tend to perform relatively the same (section 7.8), but additional considerations—such as robustness to outliers, fast hyperparameter optimization, and ability to handle large or variable step size—distinguish a few methods for specialized use (Figure 57). In some cases, a method may even be preferred for narrative reasons, having theoretical structure that matches assumptions about an underlying process or context.
@@ 3 Differentiating Analytic Functions with Static Structure
Given the dramatic success of deep learning, optimized by hardware-accelerated gradient descent implemented with libraries like JAX [jax2018github] and PyTorch [pytorch], it is natural to wonder how AutoDiff fits into the larger numerical differentiation ecosystem. In a word: AutoDiff is not for differentiating data samples, but rather for sampling fixed derivative relationships. Automatic Differentiation is most commonly used in the context of machine learning, where a model, typically named $f$ for “function”, is parameterized by big vector $\boldsymbol{\theta}$, takes a batch of inputs $X$, and produces a batch of outputs.
@@
where $L$ is a loss function, minimized when model outputs equal targets; $f$ is a model with possibly multidimensional input and output; $X$ is a set of inputs to the model; $Y$ is a set of corresponding target outputs; and $\boldsymbol{\theta}$ affects how the model produces its answers, sometimes shown after a semicolon for visual separation from model inputs. The key realization of ML is that we can consider $X$ and $Y$ to be static and given, and treat the model’s parameterization, $\boldsymbol{\theta}$, as an input, thus obtaining a relationship between overall error and the parameters. If we then differentiate this relationship with respect to the parameters and evaluate at the given input data, we obtain a gradient leading toward the loss function’s fixed point, which we can use to update model parameters as:
@@ 3.1 When Automatic Differentiation is Inappropriate
To efficiently simulate the evolution of this system, we can discretize and use 1, which employs the “leapfrog” scheme, known to be stable for the second-order wave equation when we take $\Delta t$ small enough [strang-notes, kutzbook]. At each time step, $u$ is updated by adding a small integration of its second-order spatial derivatives. This means that although the initial condition of $u$ is simply analytic, it quickly devolves to something more complicated at future time steps, possibly even non-analytic if we allow arbitrary boundary conditions.
@@
4:for $t_{n}\in\{0,\Delta t,2\Delta t,...T\}$ do
5:  $\text{righthand side}=\Big[\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big]\Big|_{\text{grid points}}$ $\triangleright$ evaluate at grid points
6:  $u_{\text{next}}=2u-u_{\text{prev}}+\Delta t^{2}\cdot\text{(righthand side)}$ $\triangleright$ leapfrog update
@@ 3.2 Automatically Differentiable Physics
Several software libraries implement this architecture, including JAX-Fluids [jaxfluids-article, jaxfluids-library], PhiFlow [phiflow], JAX-FEM [jax-fem], and Exponax [exponax]. “Differentiable Physics” [differentiable-physics] refers to the passage of gradients all the way through a physics simulation, enabling seamless assimilation in ML training and inference pipelines. Note, however, that the physical derivatives per se of such simulations are still found via whichever proximal differentiation algorithm the simulator implements (Finite Difference in the case of [jaxfluids-library, phiflow]), and so the simulation itself is only as accurate as that method rather than inheriting AutoDiff’s characteristic extreme fidelity.
@@ 4 Differentiating Noiseless Simulation Data
When a function $y(x)$ ceases to be representable with few analytic terms, it is best represented numerically, by its values. Because continuous function values are infinitely dense, the only way to practically do this is with discrete samples, $y_{n}=y(x_{n}),\ n\in\{0,...N-1\}$, for a total of $N$ points over some domain, $x\in[a,b]$. In this situation, taking derivatives becomes less exact; it is actually nonsensical to evaluate the derivative of a truly discrete sample, because a point-value has no slope. Instead, we remember samples are taken from some continuous underlying function, assume a particular interpolation to fill the gaps between samples, and differentiate that interpolation [johnson-fft-notes]. Fortunately, in the absence of noise, these approximations can be made highly accurate.
@@ 4.1 Finite Difference
This methodology can be generalized [taylor-fd] to higher, $\nu^{\text{th}}$-order, derivatives and arbitrary stencils of length $S$, $s_{0},...s_{S-1}]$, including lopsided, one-sided (“forward” or “backward” difference), and non-contiguous ones; the linear inverse problem just becomes:
@@ 4.1.1 Recommendations
Finite Difference is simple and flexible. It realizes a relationship between function and derivative values that is at once easy to implement and understand, which has made it the method of choice for several meta-differentiation libraries (section 3.2). It works with arbitrary boundary conditions. It is also fast to run, $O(Nm)$, requiring only one major iteration along the data. However, Finite Difference’s error bounds only hold for smooth functions and in the absence of noise, the area of specialty for Spectral Methods, which achieve better accuracy, so Finite Difference should only be used if Fourier (section 4.2.1) or Chebyshev (section 4.2.2) do not apply.
@@ 4.2.2 Differentiation with a Polynomial Basis: Chebyshev
Figure 8: From [dedalus], optimal sampling grids, also called “collocation points”, for a few common choices of basis.
@@
Once the Chebychev transform is achieved, then, analogous to power rule for ordinary representations of polynomials as power series, we can find Chebyshev series coefficients of the derivative in $O(N)$ using a recurrence relation [brown-cheb], which can be manipulated to produce Equation 12 [spectral-derivatives], implemented in numpy as chebder [numpy-chebder]:
@@ 4.2.3 Recommendations
Fourier spectral differentiation is best used with periodic signals. By including dramatically high frequencies, the fit can get closer at discontinuities and corners, but Gibbs phenomenon will always remain [oppenheim, trefethen2000spectral]. In some cases this problem can be alleviated or mitigated by concatenating an aperiodic signal with its mirror image (even extension) or negated mirror image (odd extension) to produce a truly periodic signal or something very close to periodic, as in Figure 11. The FFT is so fast that transforming a double-length signal is only marginally slower. But this trick does not produce a perfectly periodic extension for general signals.
@@
3 offers a method with exceptional computational efficiency and accuracy while also being unconstrained by boundary conditions. However, function samples need to come from special locations or an affine transform thereof, and the method should only be used in the absence of noise (see section 5, particularly Figure 19). The Matlab package chebfun [driscoll2014chebfun] and Python packages spectral-derivatives [spectral-derivatives] and dedalus [dedalus] exploit Chebychev polynomials to allow for accurate and efficient derivative estimates of arbitrary smooth functions.
@@ 4.3 Finite Elements
Building from piecewise elements has great benefits for flexible representation of complicated functions on irregular domains, such as shown in Figure 14. Whereas ordinary Finite Difference method requires a regular grid, [^7] Technically, the Finite Difference equations in section 4.1 can be extended to the irregular sampling case, as described in section 8.1, but this is uncommon. and Spectral methods work most naturally on rectangular domains without holes, elements are suited to work in a variety of shapes, including line segments in 1D, triangles and squares in 2D, and tetrahedra and prisms in 3D.
@@
Since $y_{h}$ is composed of basis functions, we can write it as $y_{h}=\sum_{j}c_{j}\varphi_{j}(x)$, where the $\{c_{j}\}$ are coefficients. The total number of basis functions is related to the mesh size, $N$, and the order of the basis, which can drive up the number of nodes per element. For linear basis functions as in Figure 12(a), elements only have nodes at their boundaries, so $j\in\{0,...N-1\}$.
A Galerkin projection uses the same family of functions for the basis and test sets. We have actually already encountered something which can be thought of as Galkerin projection, the Fourier transform, or more specifically the Fourier series [oppenheim, fourier-series], which goes from continuous functions on a periodic interval, say $\pi,\pi]$, to discrete coefficients, using complex exponentials:
@@
The reconstruction sum $y_{h}(x)=\sum_{j=0}^{N-1}c_{j}e^{ijx}$ leaves us with $N$ unknown coefficients, so we need $N$ equations to lock them down. These we can obtain through Equation 14 by projecting against test functions with different $k$. Plugging in the ansatz, we find:
(15) $\rightarrow\int\limits_{-\pi}^{\pi}y(x)e^{-ikx}dx=\int\limits_{-\pi}^{\pi}\Big(\sum_{j=0}^{N-1}c_{j}e^{ijx}\Big)e^{-ikx}dx\quad\text{for k = 0, ... N$-$1}$
$\rightarrow\underbrace{\langle y(x),e^{ikx}\rangle}_{\text{\small$b_{k}$}}=\sum_{j=0}^{N-1}c_{j}\underbrace{\int\limits_{-\pi}^{\pi}e^{ijx}e^{-ikx}dx}_{\text{\footnotesize$\begin{cases}2\pi&j=k\\ 0&j\neq k\end{cases}$}}=2\pi c_{k}\quad\text{for k = 0, ... N$-$1}$
@@ 4.3.1 Differentiation with Finite Elements
We have now reached a curious juncture: We can compute a piecewise basis-function representation of a function we wish to differentiate, but to do so involves inner product integrals. In practice, on a computer integrals are found via quadrature rules [larson], which use samples from particular “quadrature points” to compute the integral of a polynomial interpolant. If we provide the method with samples $y_{n}$ rather than function $y$, and cannot guarantee the spacing of these points, then trapezoidal rule is the only integration rule robust to this scenario; it requires only two quadrature points, arbitrarily spaced, which fall on the mesh points at element edges. But the trapezoidal rule interpolates between points linearly, thereby assuming that $y\in V_{h}$ already.
@@
Figure 16: Examples of functions where the notion of a weak derivative is useful. The Rectified Linear Unit (ReLU) has a sharp corner at $x=0$, making it not strictly differentiable. But intuitively its derivative should look something like a Heaviside function, which in turn has a weak derivative given by the Dirac Delta. This notion is made rigorous by the machinery of “generalized” functions (also called “distributions”) [weak-derivs], members of which can be written as a functional which maps from a function to a value via an integral against a test function, essentially identical to the form in Galerkin projection.
This setup matches the Galerkin method naturally, and it comes with the added benefit, thanks to an integration by parts trick [^9] The same manipulation is famously used in the Calculus of Variations to transfer a differential operator from the variation to the Lagrangian while spawning a boundary term, thereby enabling a factorization that yields the Euler-Lagrange equation. In higher dimension, this trick may be called Green’s First Identity [langtangen], and if the solution is vector-valued it becomes Divergence Theorem. (demonstrated below), that the solution need only be (weakly) differentiable one fewer times in space than in the original PDE, thereby enabling us to describe it with a lower-order basis, which means fewer nodes (see Figure 13) and a smaller system of equations, especially helpful in higher dimensions.
@@
$\int\limits_{0}^{L}\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\varphi_{j}(x)\varphi_{k}(x)\ dx=-\int\limits_{0}^{L}\sum_{j=1}^{N-2}c_{j}(t)\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx\quad\text{for k = 1, ... N$-$2}$
$\rightarrow$ $\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi_{j}(x)\varphi_{k}(x)\ dx}_{M_{jk}}=-\sum_{j=1}^{N-2}c_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx}_{K_{jk}}\quad\text{for k = 1, ... N$-$2}$
@@
The $(N-2)\!\times\!(N-2)$ entries of $\mathbf{M}$ and $\mathbf{K}$ do not change over time. We have transformed the PDE into a system of ODEs. Initial coefficients c can be found from initial conditions and then forward-solved with a method of lines scheme and steppers like Runge-Kutta. If the PDE instead describes a steady state with unknown $u$, the derivation culminates in an algebraic equation rather than ODEs, similar to the Fourier series example in Equation 15. For a further example, see the derivation for the 2D Poisson equation in any of [langtangen, larson, dokken, uiuc-galerkin], which is also instructive for generalizing to more spatial dimensions.

### v2

@@ Abstract
Differentiation is a cornerstone of computing and data analysis in every discipline of science and engineering. Indeed, most fundamental physics laws are expressed as relationships between derivatives in space and time. However, derivatives are rarely directly measurable and must instead be computed, often from noisy, potentially corrupt data streams. There is a rich and broad literature of computational differentiation algorithms, but many impose extra constraints to work correctly, e.g. periodic boundary conditions, or are compromised in the presence of noise and corruption. It can therefore be challenging to select the method best-suited to any particular problem. Here we review a broad range of numerical methods for calculating derivatives, present important contextual considerations and choice points, compare relative advantages, and provide basic theory for each algorithm in order to assist users with the mathematical underpinnings. This serves as a practical guide to help scientists and engineers match methods to application domains. We also provide an open-source Python package, PyNumDiff, which contains a broad suite of methods for differentiating noisy data.
@@ 1 Introduction
Numerical algorithms including Finite Differences, Spectral Methods, and Finite Elements or Volumes are now mature technologies, with the latter powering commercialized (e.g. COMSOL [comsol], Ansys and Abaqus) and open-source (e.g. FEniCS, Firedrake, OpenFOAM) PDE solvers. Likewise, Automatic Differentiation (AutoDiff) dominates in Machine Learning [jax2018github]. As practitioners in the rapidly-growing field of data-driven science, we also take particular interest in derivative estimation for noisy data streams. Motivated by the definition of the derivative from calculus, beginners in numerical computation often use finite differences by default, but this turns out to be one of the worst choices for real data. We attempt to give comprehensive coverage to all these schemes, from the narrowly focused to the most general-purpose. The overall trend is that more constrictive situations allow for better performance, so we suggest the use of specialized tools if possible. Or, if faced with a dearth of simplifying assumptions, we have made it easy for the reader to skip ahead to the pertinent category.
@@ 2 Choosing a Method
Figure 2: Suggested procedure for choosing a differentiation method. Methods are boxed with blue, violet, purple, or red to give an overall impression of how computationally efficient, accurate, robust, and user-friendly they are, with blue meaning a method excels in all respects for its specialization and red meaning a solution is hopeless. Groupings in the lowermost node represent many methods, expounded in section 7. Methods and a couple of other labels are hyperlinked for quick navigation to relevant sections.
@@
A digest of these with some preferred algorithms is given in Figure 1. Figure 2 shows a more complete and detailed overview of key considerations and methods, organized as a flowchart. The decisions therein are important because different techniques make different assumptions. For example, smoothing methods assume the signal is continuously differentiable up to some order (possibly $C^{\infty}$), and the Kalman filter requires a system model. Under changing assumptions, the most suitable approach changes, so it is important to understand whether data is periodic, noisy, etc. Notably missing among these considerations is dimensionality: Because differentiation is a linear operator, all methods extend equally well to the multidimensional case, either by multivariable chain rule, by using higher-dimensional meshes and basis functions, or by sequential application along data dimensions (in any order). We therefore primarily treat the one-dimensional case in this review, without loss of generality.
@@ 2.1 Recommendations
Down the “noisy signal, smooth derivative” path, the Fourier spectral method (i.e. transform, zero out higher modes, and inverse transform) is ideal for isolating high-frequency noise (section 5), but it should only be used when signals are periodic (can be joined smoothly end-to-end), lest Gibbs phenomenon (Figure 7) corrupt the approximation. Otherwise, one should exploit a model of system dynamics and noise when possible (section 6), although system models can be challenging to derive beyond simple systems and can be of limited utility unless fairly accurate. In the naive case (section 7), sophisticated methods tend to perform similarly (section 7.8), but additional considerations—such as robustness to outliers, fast hyperparameter optimization, and ability to handle large or variable step size—distinguish a few methods for specialized use (Figure 57). In some cases, a method may even be preferred for narrative reasons, having theoretical structure that matches assumptions about an underlying process or context.
@@ 3 Differentiating Analytic Functions with Static Structure
Given the dramatic success of deep learning, optimized by hardware-accelerated gradient descent implemented with libraries like JAX [jax2018github] and PyTorch [pytorch], it is natural to wonder how AutoDiff fits into the larger numerical differentiation ecosystem. In a word: AutoDiff is not for differentiating data samples, but rather for sampling fixed derivative relationships. Automatic Differentiation is most commonly used in the context of machine learning, where a model, typically named $f$ for “function”, is parameterized by a big vector $\boldsymbol{\theta}$, takes a batch of inputs $X$, and produces a batch of outputs.
@@
where $L$ is a loss function, minimized when model outputs equal targets; $f$ is a model with possibly multidimensional input and output; $X$ is a set of inputs to the model; $Y$ is a set of corresponding target outputs; and $\boldsymbol{\theta}$ affects how the model produces its answers, sometimes shown after a semicolon for visual separation from model inputs. The key realization of ML is that we can consider $X$ and $Y$ to be static and given, and treat the model’s parameterization, $\boldsymbol{\theta}$, as an input, thus obtaining a relationship between overall error and the parameters. If we then differentiate this relationship with respect to the parameters and evaluate at the given input data, we obtain a gradient leading toward the loss function’s minimum, which we can use to update model parameters as:
@@ 3.1 When Automatic Differentiation is Inappropriate
To efficiently simulate the evolution of this system, we can discretize and use 1, which employs the “leapfrog” scheme, known to be stable for the second-order wave equation with small enough $\Delta t$ [strang-notes, kutzbook]. At each time step, $u$ is updated by accumulating small numerical integrals of its second-order spatial derivatives. This means that although the initial condition of $u$ is simply analytic, simulated $u$ rapidly acquires more terms across future time steps. Furthermore, for arbitrary initial and boundary conditions, it is possible an analytic expression for $u$ does not exist.
@@
4:for $t_{n}\in\{0,\Delta t,2\Delta t,...,T\}$ do
5:  $\text{right-hand side}=\Big[\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\Big]\Big|_{\text{grid points}}$ $\triangleright$ evaluate at grid points
6:  $u_{\text{next}}=2u-u_{\text{prev}}+\Delta t^{2}\cdot\text{(right-hand side)}$ $\triangleright$ leapfrog update
@@ 3.2 Automatically Differentiable Physics
Several software libraries implement this architecture, including JAX-Fluids [jaxfluids-article, jaxfluids-library], PhiFlow [phiflow], JAX-FEM [jax-fem], and Exponax [exponax]. “Differentiable Physics” [differentiable-physics] refers to the passage of gradients all the way through a physics simulation, enabling seamless assimilation in ML training and inference pipelines. Note, however, that the physical derivatives per se of such simulations are still found via whichever proximal differentiation algorithm the simulator implements (Finite Difference in the case of [jaxfluids-library, phiflow]), and so the simulation itself is only as accurate as that method rather than inheriting AutoDiff’s fidelity (down to numerical precision).
@@ 4 Differentiating Noiseless Simulation Data
When a function $y(x)$ ceases to be representable with a few analytic terms, it is best represented numerically, by its values. Because continuous function values are infinitely dense, the only way to practically do this is with discrete samples, $y_{n}=y(x_{n}),\ n\in\{0,...,N-1\}$, for a total of $N$ points over some domain, $x\in[a,b]$. In this situation, taking derivatives becomes less exact because isolated samples do not allow evaluation of slope in the limit as $dx\to 0$. Instead, we remember samples are taken from some continuous underlying function, assume a particular interpolation to fill the gaps between samples, and differentiate that interpolation [johnson-fft-notes]. Fortunately, in the absence of noise, these approximations can be made highly accurate.
@@ 4.1 Finite Difference
This methodology can be generalized [taylor-fd] to higher, $\nu^{\text{th}}$-order, derivatives and arbitrary stencils of length $S$, $s_{0},...,s_{S-1}]$, including lopsided, one-sided (“forward” or “backward” difference), and non-contiguous ones; the linear inverse problem just becomes:
@@ 4.1.1 Recommendations
Finite Difference is simple and flexible. It realizes a relationship between function and derivative values that is at once easy to implement and understand, which has made it the method of choice for several meta-differentiation libraries (section 3.2). It works with arbitrary boundary conditions and is fast to run, $O(Nm)$, requiring only one major iteration along the data. However, Finite Difference’s error bounds only hold for smooth functions and in the absence of noise, the same situation where Spectral Methods, which achieve better accuracy, can excel, so Finite Difference should only be used if Fourier (section 4.2.1) or Chebyshev (section 4.2.2) do not apply.
@@ 4.2.2 Differentiation with a Polynomial Basis: Chebyshev
Figure 8: From [dedalus], optimal sampling grids across interval domains, also called “collocation points”, for a few common choices of basis. Each basis is defined on its own canonical interval, e.g. $0,2\pi)$ for Fourier and [1,1][-1,1] for Chebyshev, so other intervals are affine-transformed before basis function fit, which is easily reversible to return to the original domain.
@@
Once the Chebyshev transform is achieved, then, analogous to the power rule for ordinary representations of polynomials as power series, we can find Chebyshev series coefficients of the derivative in $O(N)$ using a recurrence relation [brown-cheb], which can be manipulated to produce Equation 12 [spectral-derivatives], implemented in numpy as chebder [numpy-chebder]:
@@ 4.2.3 Recommendations
Fourier spectral differentiation is best used with periodic signals. By including higher frequencies, the fit can get closer at discontinuities and corners, but Gibbs phenomenon will always remain [oppenheim, trefethen2000spectral]. In some cases this problem can be alleviated or mitigated by concatenating an aperiodic signal with its mirror image (even extension) or negated mirror image (odd extension) to produce a truly periodic signal or something very close to periodic, as in Figure 11. The FFT is so fast that transforming a double-length signal is only marginally slower. But this trick does not produce a perfectly periodic extension for general signals.
@@
3 offers a method with exceptional computational efficiency and accuracy while also being unconstrained by boundary conditions. However, function samples need to come from special locations or an affine transform thereof, and the method should only be used in the absence of noise (see section 5, particularly Figure 19). The Matlab package chebfun [driscoll2014chebfun] and Python packages spectral-derivatives [spectral-derivatives] and dedalus [dedalus] exploit Chebyshev polynomials to allow for accurate and efficient derivative estimates of arbitrary smooth functions.
@@ 4.3 Finite Elements
Building from piecewise elements has great benefits for flexible representation of complicated functions on irregular domains, such as shown in Figure 14. Whereas ordinary Finite Difference requires a regular grid, [^7] Technically, the Finite Difference equations in section 4.1 can be extended to the irregular sampling case, as described in section 8.1, but this is uncommon. and Spectral Methods work most naturally on rectangular domains without holes, elements are suited to work in a variety of shapes, including line segments in 1D, triangles and squares in 2D, and tetrahedra and prisms in 3D.
@@
Since $y_{h}$ is composed of basis functions, we can write it as $y_{h}=\sum_{j}c_{j}\varphi_{j}(x)$, where the $\{c_{j}\}$ are coefficients. The total number of basis functions is related to the mesh size, $N$, and the order of the basis, which can drive up the number of nodes per element. For linear basis functions as in Figure 12(a), elements only have nodes at their boundaries, so $j\in\{0,...,N-1\}$.
A Galerkin projection uses the same family of functions for the basis and test sets. We have actually already encountered something which can be thought of as Galerkin projection, the Fourier transform, or more specifically the Fourier series [oppenheim, fourier-series], which maps from a continuous function on a periodic interval, say $\pi,\pi]$, to discrete coefficients, using complex exponentials:
@@
The reconstruction sum $y_{h}(x)=\sum_{j=0}^{N-1}c_{j}e^{ijx}$ leaves us with $N$ unknown coefficients, so we need $N$ equations to lock them down. These we can obtain through Equation 14 by projecting against test functions with different $k$. Plugging in the ansatz for $y_{h}$, we find:
(15) $\rightarrow\int\limits_{-\pi}^{\pi}y(x)e^{-ikx}dx=\int\limits_{-\pi}^{\pi}\Big(\sum_{j=0}^{N-1}c_{j}e^{ijx}\Big)e^{-ikx}dx\quad\text{for }k=0,.,N-1$
$\rightarrow\underbrace{\langle y(x),e^{ikx}\rangle}_{\text{\small$b_{k}$}}=\sum_{j=0}^{N-1}c_{j}\underbrace{\int\limits_{-\pi}^{\pi}e^{ijx}e^{-ikx}dx}_{\text{\footnotesize$\begin{cases}2\pi&j=k\\ 0&j\neq k\end{cases}$}}=2\pi c_{k}\quad\text{for }k=0,.,N-1$
@@ 4.3.1 Differentiation with Finite Elements
We have now reached a curious juncture: We can compute a piecewise basis-function representation of a function we wish to differentiate, but to do so involves inner product integrals. In practice, on a computer integrals are found via quadrature rules [larson], which use samples from particular “quadrature points” to compute the integral of a polynomial interpolant. If we provide the method with samples $y_{n}$ rather than function $y$, and cannot guarantee the spacing of these points, then trapezoidal rule (or its higher-dimensional analog) is the only integration rule robust to this scenario, requiring only quadrature points at element boundaries. But the trapezoidal rule interpolates between vertices linearly, thereby assuming that $y\in V_{h}$ already.
@@
Figure 16: Examples of functions where the notion of a weak derivative is useful. The Rectified Linear Unit (ReLU) has a sharp corner at $x=0$, making it not strictly differentiable. But intuitively its derivative should look something like a Heaviside function, which in turn has a weak derivative given by the Dirac Delta. This notion is made rigorous by the machinery of “generalized” functions (also called “distributions”) [weak-derivs], members of which can be written as a functional which maps from function to value via an integral against a test function, essentially identical to the form in Galerkin projection.
This setup matches the Galerkin method naturally, and it comes with the added benefit, thanks to an integration by parts trick [^9] The same manipulation is famously used in Calculus of Variations to transfer a differential operator from the variation to the Lagrangian while spawning a boundary term, thereby enabling a factorization that yields the Euler-Lagrange equation. In higher dimension, this trick may be called Green’s First Identity [langtangen], a special case of Divergence Theorem. (demonstrated below), that the solution need only be (weakly) differentiable one fewer times in space than in the original PDE, thereby enabling us to describe it with a lower-order basis, which means fewer nodes (see Figure 13) and a smaller system of equations, especially helpful for saving computation in higher dimensions. [^10] The weak formulation may not be strong, but it is powerful.
@@
$\int\limits_{0}^{L}\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\varphi_{j}(x)\varphi_{k}(x)\ dx=-\int\limits_{0}^{L}\sum_{j=1}^{N-2}c_{j}(t)\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx\quad\text{for }k=1,...,N-2$
$\rightarrow$ $\sum_{j=1}^{N-2}\ddot{c}_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi_{j}(x)\varphi_{k}(x)\ dx}_{M_{jk}}=-\sum_{j=1}^{N-2}c_{j}(t)\underbrace{\int\limits_{0}^{L}\varphi^{\prime}_{j}(x)\varphi^{\prime}_{k}(x)\ dx}_{K_{jk}}\quad\text{for }k=1,...,N-2$
@@
The $(N-2)\!\times\!(N-2)$ entries of $\mathbf{M}$ and $\mathbf{K}$ do not change over time. We have transformed the PDE into a system of ODEs. Initial coefficients c can be found from initial conditions and then forward-solved with a method of lines scheme and steppers like Runge-Kutta. If the PDE instead describes a steady state with unknown $u$, the derivation culminates in an algebraic equation rather than ODEs, similar to the Fourier series example in Equation 15. For a further example, see the derivation for the 2D Poisson equation in any of [langtangen, larson, dokken, olver, uiuc-galerkin], which is also instructive for generalizing to more spatial dimensions.